in 3years time george's father will be as old as george if georges father is now 37 years,how old is george now.
2x+3+x+3=37
2x+x=37-3-3
3x/3=37/3
x=12.33

Found 3 solutions by ikleyn, greenestamps, Theo:
Answer by ikleyn(52794)

.

in 3years time george's father will be as old as george if georges father is now 37 years,how old is george now.
2x+3+x+3=37
2x+x=37-3-3
3x/3=37/3
x=12.33
~~~~~~~~~~~~~~~~~~~

                    Nonsense.


    It is really a rare case to meet a nonsense of such extent.




Answer by greenestamps(13200)



this is nonsense....

With humans at least, a father will never be the same age as his son.

read what you wrote; correct it, and post the problem correctly

Answer by Theo(13342)

your problem statement and your equations do not track with each other.

the problem statement doesn't make any sense.

in 3 years time george's father will be as old as george is now?

that makes george older than his father.

not likely.

your equation suggests that, in 3 years time, george's father will be twice as old as george.

that makes more sense.

assuming that's what you mean, then:

in 3 years time, george's father will be twice as old as george.

if george's father is now 37, how old is george.

let x = george's age now.

in 3 years time, george will be x + 3 years old.

in 3 years time, george's father will be 37 + 3 years = 40 years old.

your equation is 40 = 2 * (x + 3)

simplify to get 40 = 2x + 6

subtract 6 from both sides of this equation to get 34 = 2x

solve for x to get x = 34/2 = 17

george's father is now 37 and george is 17.

in 3 years, george's father sill be 40 and george will be 20.

in 3 years timw, george's father is 40 / 20 = twica as old as george.

if you meant:

in 3 years time george's father will be twice as old as george is now.

then the equation would be:

40 = 2x and you will get that george is 20 now.